A large chunk of ice with mass of 15.0kg falls from a roof 8.0m above the ground. Assume air resistance is negligible, kinetic energy is 14.25 J. Determine the speed of the ice when it reaches the ground.


1
Expert's answer
2022-01-19T12:40:19-0500

Apply conservation of energy:


"\dfrac12mv^2=KE_i+mgh,\\\space\\
v=\sqrt{2(KE_i/m+gh)}=\sqrt{2(14.25/15+9.8·8)},\\
v=13.6\text{ m/s}."